Job 605 has a total cost of $35,820. The MOH Rate is $0.75 per direct labor...

  1. Job 605 has a total cost of $35,820. The MOH Rate is $0.75 per direct labor cost (i.e. 75% of direct labor cost). $11,830of overhead has been allocated to the job. What was the amount of direct materials traced to the job?

    8,217

    15,118

    20,047

    23,990
Answer #1

ANSWER=8217

WORKING

DIRECT LABOUR=MANUFACTURING OVERHEAD/75%

= 11830/75%=15773

TOTAL COST OF JOB 605=DIRECT MATERIAL+DIRECT LABOUR +MANUFACTURING OVERHEAD

35820=DIRECT MATERIAL+15773+11830

DIRECT MATERIAL=35820-15773-11830=8217